# The Algorithmic Gaze: An Audit and Ethnography of the LAION-Aesthetics Predictor Model

### v2

@@ Abstract.
Visual generative AI models are trained using a one-size-fits-all measure of aesthetic appeal. However, what is deemed “aesthetic” is inextricably linked to personal taste and cultural values, raising the question of whose taste is represented in visual generative AI models. In this work, we study an aesthetic evaluation model—LAION Aesthetic Predictor (LAP)—that is widely used to curate datasets to train visual generative image models, like Stable Diffusion, and evaluate the quality of AI-generated images. To understand what LAP measures, we audited the model across three datasets. First, we examined the impact of aesthetic filtering on the LAION-Aesthetics Dataset ($\sim$1.2B images), which was curated from LAION-5B using LAP. We find that the LAP disproportionally filters in images with captions mentioning women, while filtering out images with captions mentioning men or LGBTQ+ people. Then, we used LAP to score $\sim$330k images across two art datasets, finding the model rates realistic images of landscapes, cityscapes, and portraits from western and Japanese artists most highly. In doing so, the algorithmic gaze of this aesthetic evaluation model reinforces the imperial and male gazes found within western art history. In order to understand where these biases may have originated, we performed a digital ethnography of public materials related to the creation of LAP. We find that the development of LAP reflects the biases we found in our audits, such as the aesthetic scores used to train LAP primarily coming from English-speaking photographers and western AI-enthusiasts. In response, we discuss how aesthetic evaluation can perpetuate representational harms and call on AI developers to shift away from prescriptive measures of ”aesthetics” toward more pluralistic evaluation.
AI, Art, Aesthetic Evaluation
copyright: acmcopyrightjournalyear: 2025ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Although artistic taste is deeply subjective and highly political (Bourdieu, 1984), AI researchers often attempt to quantify the aesthetic quality of media (Tjandra et al., 2025; Huang et al., 2024). While aesthetic quality assessment (AQA) has been a topic of research in the computer vision community for decades (Datta et al., 2006; Murray et al., 2012), this task is particularly important to examine in light of the rapid growth in visual generative AI. Not only are AQA models used to curate the training data for visual generative AI models (Chen et al., 2024), but they are also used to assess the quality of AI-generated images (Gao and Liu, 2024). As generative AI models are trained on datasets scraped from across the web, researchers rely on computational methods to filter out “low-quality” data. Prior work has called attention to the myriad gender and cultural biases in common dataset filtering techniques—such as blocklists (Dodge et al., 2021) and image-text similarity metrics (Hong et al., 2024)—but the consequences of aesthetic filtering has yet to be explored in detail.
@@
Due to the subjectivity of aesthetic evaluation, we first explored (RQ1) what the LAP model determines to be a “high-quality” image using three datasets. First, we examined the impact of aesthetic filtering on the LAION-Aesthetics Dataset ($\sim$1.2B images), which was curated using LAP. We find that the LAP disproportionally filters in images with captions mentioning women, while filtering out images with captions mentioning men or LGBTQ+ people. Then, we used LAP to score $\sim$330k images across two art datasets, finding the model rates realistic images of landscapes, cityscapes, and portraits from western and Japanese artists most highly. In other words, the algorithmic gaze of LAP reproduces the male gaze within western art history (Berger, 2008). In order to understand where these biases may have originated, we performed a digital ethnography of public materials to examine (RQ2) how was LAP developed. We find that the development of LAP reflects the biases we found in our audits, such as the aesthetic scores used to train LAP primarily coming from English-speaking photographers and western AI-enthusiasts.
@@ 2.1. Background on LAION-Aesthetics
LAION has faced substantial criticism from both AI researchers and artists. LAION researchers compiled the 5.86 billion image-text pairs of the LAION-5B dtaset by extracting image links and alternative text descriptions from the IMG HTML tags of websites archived by the Common Crawl, a US-based non-profit that has scraped publicly available websites since 2008 (Reisner, 2025; Baack, 2024b). Due to LAION researchers’ indiscriminate data collection practices, researchers found that LAION-5B contained over a thousand images of child sexual abuse material (CSAM) (Thiel, 2023). In addition to containing CSAM, LAION-5B contains millions of copyrighted images, leading to numerous lawsuits against LAION (Filikhina and Elsaß, 2025) and Stability AI (Tobin, 2025). Furthermore, Birhane et al. found that the image descriptions in LAION-400M and the English language subset of LAION-5B (LAION-2B-en) often contain hate speech (Birhane et al., 2023).
@@
Although prior research has examined biases in LAION datasets (Birhane et al., 2023; Hong et al., 2024), the LAION-Aesthetics Dataset and the LAION-Aesthetics Predictor (LAP) model used to curate this dataset are underexplored. Not only is the LAION-Aesthetics Dataset widely used to pre-train and fine-tune visual generative AI models, but the LAP model used to to curate this dataset has also been used to curate internal corporate datasets (Chen et al., 2024) and to evaluate the performance of visual generative AI models (Ding et al., 2024; Fong and See, 2024; Gao and Liu, 2024; He et al., 2023). Although some humanistic research has critiqued how LAP was trained (Itävuori, 2025; Vianna, 2025), researchers have not yet empirically investigated what this image quality assessment model measures in practice. In doing so, we contribute both to scholarship on art and AI as well as, more broadly, AI evaluation and alignment practices. In the following sections, we briefly describe each of these contexts in greater detail.
2.2. AI Evaluation & Alignment
@@
AI researchers are increasingly critiquing the measurements used to train and evaluate AI models. For example, benchmarks often claim to measure concepts with no agreed upon definition, such as “intelligence” (Bean et al., 2025). Researchers have also criticized the notion of universal “gold standards” in data annotation (Wang et al., 2024). While researchers often treat annotator disagreement as noise to be filtered out, disagreements can stem from annotators’ differing lived experiences (Wang et al., 2024; Sap et al., 2022). In response, researchers have called for distinguishing between datasets intended to represent different beliefs or encode one specific belief (Röttger et al., 2022). AI alignment research focuses on normative evaluations where people can reasonably disagree about how a model should behave. For instance, LLMs tend to reflect social values more commonly held in western countries according to national value surveys (Tao et al., 2024). At the same time, methodology can flatten intra-cultural differences (AlKhamissi et al., 2025). Instead of choosing one set of values to reflect, researchers have increasingly called for designing AI systems more pluralistically (Sorensen et al., 2024a, b). In this work, we contend with “aesthetics” as an underexamined axis of alignment.
@@
Prior work has found that design of visual generative AI models often harms marginalized communities. Text-to-image (T2I) models often stereotype or poorly represent women (Bianchi et al., 2023), LGBTQ+ people (Ungless et al., 2023; Taylor et al., 2025), people of color (Wolfe and Caliskan, 2022), and non-western cultures (Fu et al., 2024; Mim et al., 2024; Qadri et al., 2025; Yerukola et al., 2025). T2I models made by western tech companies tend to perform worse in languages other than English, such as Farsi (Qadri et al., 2025) or Bangla (Mim et al., 2024). This linguistic bias can also be seen in a dataset we examine in this study: the LAION-Aesthetics Dataset is restricted to images with English captions (Schuhmann, 2022). HCI researchers have found that artists often enjoy the glitches and strangeness of visual generative AI models (Chang et al., 2023), leading some to critique the stylistic bias toward realism in newer models (Taylor et al., 2025). In light of this concern, Taylor et al. have called for future research “examining measurements of aesthetic quality in popular datasets” (Taylor et al., 2025), which we take up in this work.
@@
While the artistic consequences of aesthetic evaluation are underexamined, aesthetic quality assessment (AQA) has been a subject of research in the computer vision community for decades (Murray et al., 2012; Datta et al., 2006). In this work, researchers formulate AQA as a regression problem predicting a numerical aesthetic score. At the same time, this de-contextualized measurement of aesthetic quality has been critiqued (Goree et al., 2025; Huang et al., 2024). Goree et al. (2025) highlight value misalignments between various AQA models and human-preferences through the development of an aesthetic evaluation mobile camera application. Like ourselves, Li et al. (2025) recently advocated for investigating aesthetic biases as an underexamined avenue of representational harms in mulitmodal LLMS. However, the authors conceptualize aesthetic biases as a distributional alignment problem wherein aesthetic evaluation models may privilege the values of one demographic group over others, specially studying age, gender and education. We extend this work by examining the broader cultural consequences of AQA on data curation and model evaluation, grounding the task within art history.
@@ 3.1. Audit Methodology
We audited LAP using three datasets. First, we examine the LAION-Aesthetics Dataset that LAP was designed to curate from LAION-5B. We then audited LAP using two art datasets with robust metadata on the artists, cultures, styles, and mediums associated with each image. The first dataset consists of 249,351 public-domain images from the Metropolitan Museum of Art (MET).  [^2] https://www.metmuseum.org/hubs/open-access Located in New York City, the MET is the fourth largest museum in the world and its collection spans over 5,000 years (Tapia, 2021). While the MET dataset contains images from a wide breadth of art across various cultures and time periods, few images are of modern or contemporary art (only 865 images). In response, we decided to evaluate LAP on the WikiArt dataset (Saleh and Elgammal, 2015) of 81,444 images of visual art by 129 primarily modern (mid-1800s to mid-1900s) artists. WikiArt is a community edited visual art encyclopedia commonly studied in digital humanities research (Mohamed et al., 2022; Fu and Razmjooy, 2025).
Across each datasets, our analyses focus on understanding differences between images rated above and below an aesthetics score threshold of 6.5. We chose this threshold because computer vision researchers often use images scored above 6.5 by LAP to represent high-quality images. The LAION-Aesthetics Dataset was released in multiple subsets based on different aesthetic thresholds, with images scored 4.5+ ($\sim$ 1.2B images) being the lowest rated dataset and 6.5+ ($\sim$ 625k images) being the highest rated. The LAD 6.5+ is widely used by computer vision researchers to fine-tune generative image models (Guo et al., 2024; Castells et al., 2024), evaluate new training methods (Luo et al., 2023; Fan et al., 2024), and create new datasets (Tudosiu et al., 2024; Li et al., 2024a). Additionally, Ding et al. measure the effectiveness of data poisoning attacks on text-to-image models by defining ”generation aesthetics” as the ”% of images with [LAP] aesthetics ¿ 6.5” (Ding et al., 2024).
@@ 3.2. Findings: The LAION-Aesthetics Dataset
Figure 1. Top 25 domains of images in the LAION-Aesthetics Dataset rated 6.5+ by the LAION-Aesthetic Prediction model
Figure 2. Pointwise Mutual Information (PMI) between regex and images being included in the 6.5+ subset of the LAION-Aesthetics Dataset. Terms with higher PMI (e.g., wom[ae]n) have higher likelihood of being included.
@@
We find substantial differences between the URLs of images below and above the 6.5 aesthetic threshold. In LAD 4.5+, the top 25 most common domains account for about 36% of all the images in the dataset. These domains are primarily from e-commerce websites (e.g., Shopify), stock image sites (e.g., Getty Images), and web hosts (e.g., Wordpress). For LAD 6.5+, the top 25 image domains account for proximately 52% of the dataset, meaning a majority of the images in this dataset come from the handful of websites in Figure 1. Of these 25 domains, 13 were also in the top 25 domains for LAD 4.5+. However, the other domains are primarily from smaller websites where independent visual artists (e.g., Redbubble, DeviantArt) and photographers (e.g., SmugMug, Flikr, 500px) go to share or sell their work. Additionally, the 25th most common source of images in LAD 6.5 are from a small nature photography agency, robertharding.com. This indicates that LAP may rate photographic images more highly.
We find that images with descriptions mentioning women are more likely to be included in LAD 6.5+, while images mentioning men or LGBTQ+ communities are more likely to be excluded. This finding echoes scholarship on the heteronormative, male gaze in western art (Mulvey, 1975; Berger, 2008). According to Berger (2008) “Women [in western art] are depicted in a quite different way from men — not because the feminine is different from the masculine — but because the ideal spectator is always assumed to be male and the image of the woman is designed to flatter him.”
We also find that images with descriptions mentioning Hindu, Buddhist, or Christian communities are more likely to be included in LAD 6.5+, while images with descriptions mentioning Jews or Muslims are more likely to be filtered out. Additionally, images mentioning Asian or African communities are more likely to be included in LAD 6.5+ than those mentioning Latinx/e, Caucasian, or European communities. This may suggest that LAP more highly rates images of Asian and African communities than European communities. Alternatively, images containing Caucasians or Europeans may not mention this detail in their captions because whiteness is an unmarked racial category (Brekhus, 1998). Understanding the cultural implications of aesthetic evaluation requires considering more than the representation of identity groups (Taylor et al., 2025). Below, we explore aesthetic filtering in greater detail by examining how LAP rates art from various cultures.
@@
European Paintings 5.731 35 2,289
Photographs 5.405 15 6,685
Drawings & Prints 5.063 13 60,889
Robert Lehman Collection 5.035 3 2,273
The Cloisters 4.964 0 2,258
Costume Institute 4.941 0 8,124
Asian Art 4.931 60 30,597
The American Wing 4.856 49 12,292
Arms & Armor 4.823 0 7,260
The Libraries 4.772 0 107
Musical Instruments 4.750 0 2,498
Modern & Contemporary Art 4.731 1 865
European Sculpture & Decorative Arts 4.717 1 37,709
@@
Medieval Art 4.697 0 6,887
Egyptian Art 4.647 0 13,269
Islamic Art 4.555 0 13,054
Ancient West Asian Art 4.414 0 6,093
Greek & Roman Art 4.295 0 30,011
@@
3.3. Findings: The Metropolitan Museum of Art Dataset

### v3

# The Algorithmic Gaze of Image Quality Assessment: An Audit and Trace Ethnography of the LAION-Aesthetics Predictor

@@ Abstract.
Visual generative AI models are trained using a one-size-fits-all measure of aesthetic appeal. However, what is deemed “aesthetic” is inextricably linked to personal taste and cultural values, raising the question of whose taste is represented in visual generative AI models. In this work, we study an aesthetic evaluation model—the LAION-Aesthetics Predictor (LAP)—that is widely used to curate datasets to train visual generative image models, like Stable Diffusion, and evaluate the quality of AI-generated images. To understand what LAP measures, we audited the model across three datasets. First, we examined the impact of aesthetic filtering on the LAION-Aesthetics Dataset ($\sim$1.2B images), which was curated from LAION-5B using LAP. We find that the LAP disproportionally filters in images with captions mentioning women, while filtering out images with captions mentioning men or LGBTQ+ people. Then, we used LAP to score $\sim$330k images across two art datasets, finding the model rates realistic images of landscapes, cityscapes, and portraits from western and Japanese artists most highly. In doing so, the algorithmic gaze of this aesthetic evaluation model reinforces the imperial and male gazes found within western art history. In order to understand where these biases may have originated, we performed a trace ethnography of public materials related to the creation of LAP. We find that the development of LAP reflects the biases we found in our audits, such as the aesthetic scores used to train LAP primarily coming from English-speaking photographers and western AI-enthusiasts. In response, we discuss how aesthetic evaluation can perpetuate representational harms and call on AI developers to shift away from prescriptive measures of “aesthetics” toward more pluralistic evaluation.
AI, Art, Audit, Alignment, Trace Ethnography, Aesthetic Evaluation, Image Quality Assessment, Aesthetic Quality Assessment
copyright: acmcopyrightjournalyear: 2025conference: FAccT ’26; June 25–28, 2026; Montréal, Canadaccs: Computing methodologies Computer visionccs: Human-centered computing Empirical studies in HCI* Co-senior authors contributed equally to this research
@@ 1. Introduction
Although artistic taste is deeply subjective and highly political (Bourdieu, 1984), AI researchers often attempt to quantify the aesthetic quality of media (Tjandra et al., 2025; Huang et al., 2024). While aesthetic quality assessment (AQA) has been a topic of research in the computer vision community for decades (Datta et al., 2006; Murray et al., 2012), this task is particularly important to examine in light of the rapid growth in visual generative AI. Not only are AQA models used to curate the training data for visual generative AI models (Chen et al., 2024), but they are also used to assess the quality of AI-generated images (Gao and Liu, 2024). As generative AI models are trained on datasets scraped from across the web, researchers rely on computational methods to filter out “low-quality” data. Prior work has called attention to myriad gender and cultural biases in common dataset filtering techniques—such as blocklists (Dodge et al., 2021) and image-text similarity metrics (Hong et al., 2024)—but the consequences of aesthetic filtering has yet to be explored in detail.
@@
Due to the subjectivity of aesthetic evaluation, we first explored (RQ1) what the LAP model determines to be a “high-quality” image using three datasets. First, we examined the impact of aesthetic filtering on the LAION-Aesthetics Dataset ($\sim$1.2B images), which was curated using LAP. We find that the LAP disproportionally filters in images with captions mentioning women, while filtering out images with captions mentioning men or LGBTQ+ people. Then, we used LAP to score $\sim$330k images across two art datasets, finding the model rates realistic images of landscapes, cityscapes, and portraits from western and Japanese artists most highly. In other words, the algorithmic gaze of LAP reproduces the male gaze within western art history (Berger, 2008). In order to understand where these biases may have originated, we performed a trace ethnography of public materials to examine (RQ2) how LAP was developed. We find that the development of LAP reflects the biases we found in our audits, such as the aesthetic scores used to train LAP primarily coming from English-speaking photographers and western AI-enthusiasts.
@@ 2.1. Background on LAION-Aesthetics
LAION has faced substantial criticism from both AI researchers and artists. LAION researchers compiled the 5.86 billion image-text pairs of the LAION-5B dataset by extracting image links and alternative text descriptions from the IMG HTML tags of websites archived by the Common Crawl, a US-based non-profit that has scraped publicly available websites since 2008 (Reisner, 2025; Baack, 2024b). Due to LAION researchers’ indiscriminate data collection practices, researchers found that LAION-5B contained over a thousand images of child sexual abuse material (CSAM) (Thiel, 2023). In addition to containing CSAM, LAION-5B contains millions of copyrighted images, leading to numerous lawsuits against LAION (Filikhina and Elsaß, 2025) and Stability AI (Tobin, 2025). Furthermore, Birhane et al. found that the image descriptions in LAION-400M and the English language subset of LAION-5B (LAION-2B-en) often contain hate speech (Birhane et al., 2023).
@@
Although prior research has examined biases in LAION datasets (Birhane et al., 2023; Hong et al., 2024), the LAION-Aesthetics Dataset and the LAION-Aesthetics Predictor (LAP) model [^2] https://github.com/christophschuhmann/improved-aesthetic-predictor used to curate this dataset are underexplored. Not only is the LAION-Aesthetics Dataset widely used to pre-train and fine-tune visual generative AI models, but the LAP model used to to curate this dataset has also been used to curate internal corporate datasets (Chen et al., 2024) and to evaluate the performance of visual generative AI models (Ding et al., 2024; Fong and See, 2024; Gao and Liu, 2024; He et al., 2023). Although some humanistic research has critiqued how LAP was trained (Itävuori, 2025; Vianna, 2025), researchers have not yet empirically investigated what this image quality assessment model measures in practice. In doing so, we contribute both to scholarship on art and AI as well as, more broadly, AI evaluation and alignment practices. In the following sections, we briefly describe each of these contexts in greater detail.
$2.$ AI Evaluation & Alignment
@@
AI researchers are increasingly critiquing the measurements used to train and evaluate AI models. For example, benchmarks often claim to measure concepts with no agreed upon definition, such as “intelligence” (Bean et al., 2025). Researchers have also criticized the notion of universal “gold standards” in data annotation (Wang et al., 2024). While researchers often treat annotator disagreement as noise to be filtered out, disagreements can stem from annotators’ differing lived experiences (Wang et al., 2024; Sap et al., 2022). In response, researchers have called for distinguishing between datasets intended to represent different beliefs or encode one specific belief (Röttger et al., 2022). AI alignment research focuses on normative evaluations where people can reasonably disagree about how a model should behave. For instance, LLMs tend to reflect social values more commonly held in western countries according to national value surveys (Tao et al., 2024). At the same time, this methodology can flatten intra-cultural differences (AlKhamissi et al., 2025). Instead of choosing one set of values to reflect, researchers have increasingly called for designing AI systems more pluralistically (Sorensen et al., 2024a, b). In this work, we contend with “aesthetics” as an underexamined axis of alignment.
@@
Prior work has found that the design of visual generative AI models often harms marginalized communities. Text-to-image (T2I) models stereotype or poorly represent women (Bianchi et al., 2023), LGBTQ+ people (Ungless et al., 2023; Taylor et al., 2025), people of color (Wolfe and Caliskan, 2022), and non-western cultures (Fu et al., 2024; Mim et al., 2024; Qadri et al., 2025; Yerukola et al., 2025). T2I models made by western tech companies tend to perform worse in languages other than English, such as Farsi (Qadri et al., 2025) or Bangla (Mim et al., 2024). This linguistic bias can also be seen in a dataset we examine in this study: the LAION-Aesthetics Dataset is restricted to images with English captions (Schuhmann, 2022). HCI researchers have found that artists often enjoy the glitches and strangeness of T2I models (Chang et al., 2023), leading some to critique the stylistic bias toward realism in newer models (Taylor et al., 2025). In light of this concern, Taylor et al. have called for future research “examining measurements of aesthetic quality in popular datasets” (Taylor et al., 2025), which we take up in this work.
@@
While the artistic consequences of aesthetic evaluation are underexamined, aesthetic quality assessment (AQA) has been a subject of research in the computer vision community for decades (Murray et al., 2012; Datta et al., 2006). In this work, researchers formulate AQA as a regression problem predicting a numerical aesthetic score. At the same time, this de-contextualized measurement of aesthetic quality has been critiqued (Goree et al., 2025; Huang et al., 2024). Goree et al. (2025) highlight value misalignments between various AQA models and human-preferences by developing an aesthetic evaluation mobile camera application. Like ourselves, Li et al. (2025) recently advocated for investigating aesthetic biases as an underexamined form of representational harms in mulit-modal LLMs. However, the authors conceptualize aesthetic biases as a distributional alignment problem wherein aesthetic evaluation models may privilege the values of one demographic group over others, examining age, gender and education. We extend this work by studying the broader cultural consequences of AQA on data curation and model evaluation, grounding the task within art history.
@@ 3.1. Audit Methodology
We audited LAP using three datasets. First, we examine the LAION-Aesthetics Dataset curated from LAION-5B by LAP. We then audited LAP using two art datasets with robust metadata on the artists, cultures, styles, and mediums associated with each image. The first dataset consists of 249,351 public-domain images from the Metropolitan Museum of Art (MET). [^3] https://www.metmuseum.org/hubs/open-access Located in New York City, the MET is the fourth largest museum in the world and its collection spans over 5,000 years (Tapia, 2021). While the MET dataset contains images from a wide breadth of art across various cultures and time periods, few images are of modern or contemporary art (only 865 images). Therefore, we evaluated LAP on the WikiArt dataset (Saleh and Elgammal, 2015) of 81,444 images of visual art by 129 primarily modern (mid-1800s to mid-1900s) artists. WikiArt is a community edited visual art encyclopedia commonly studied in digital humanities research (Mohamed et al., 2022; Fu and Razmjooy, 2025).
Across each datasets, our analyses primarily focus on understanding differences between images rated above and below an aesthetics score of 6.5. We chose this threshold because it is commonly used in computer vision research to represent high-quality images. The LAION-Aesthetics Dataset was released in multiple subsets based on different aesthetic thresholds, with images scored 4.5+ ($\sim$ 1.2B images) being the lowest rated dataset and 6.5+ ($\sim$ 625k images) being the highest rated. The LAD 6.5+ is widely used by computer vision researchers to fine-tune generative image models (Guo et al., 2024; Castells et al., 2024), evaluate new training methods (Luo et al., 2023; Fan et al., 2024), and create new datasets (Tudosiu et al., 2024; Li et al., 2024a). Additionally, Ding et al. measure the effectiveness of data poisoning attacks on text-to-image models by defining “generation aesthetics” as the “% of images with [LAP] aesthetics ¿ 6.5” (Ding et al., 2024). In Tables 1 and 2 we also show the number of images rated 6+ to demonstrate that the aesthetic biases we identify across the 6.5+ threshold can also be seen across the 6+ threshold.
@@ 3.2. Findings: The LAION-Aesthetics Dataset
Figure 1. Top 25 domains of images in the LAION-Aesthetics Dataset rated 6.5+ by the LAION-Aesthetics Prediction model
[…]
@@
We find substantial differences between the URLs of images below and above the 6.5 aesthetic threshold. In LAD 4.5+, the top 25 most common domains account for about 36% of all the images in the dataset. These domains are primarily from e-commerce websites (e.g., Shopify), stock image sites (e.g., Getty Images), and web hosts (e.g., Wordpress). For LAD 6.5+, the top 25 image domains account for proximately 52% of the dataset, meaning a majority of the images in this dataset come from the handful of websites in Figure 1. Of these 25 domains, 13 were also in the top 25 domains for LAD 4.5+. However, the other domains are primarily from smaller websites where independent visual artists (e.g., Redbubble, DeviantArt) and photographers (e.g., SmugMug, Flikr, 500px) share or sell their work. Additionally, the 25th most common source of images in LAD 6.5+ are from a small nature photography agency, robertharding.com. This indicates that LAP may rate photographic images more highly.
@@
Scored $\geq$ 6
Num Images
European Paintings 5.73 35 483 2,289
Photographs 5.40 15 548 6,685
Drawings & Prints 5.06 13 435 60,889
Robert Lehman Collection 5.04 3 41 2,273
The Cloisters 4.96 0 1 2,258
Costume Institute 4.94 0 7 8,124
Asian Art 4.93 60 1,240 30,597
The American Wing 4.86 49 474 12,292
Arms & Armor 4.82 0 2 7,260
The Libraries 4.77 0 0 107
Musical Instruments 4.75 0 3 2,498
Modern & Contemporary Art 4.73 1 11 865
European Sculpture & Decorative Arts 4.72 1 12 37,709
@@
Medieval Art 4.70 0 1 6,887
Egyptian Art 4.65 0 3 13,269
Islamic Art 4.56 0 15 13,054
Ancient West Asian Art 4.41 0 1 6,093
Greek & Roman Art 4.29 0 7 30,011
@@
We find that images with descriptions mentioning women are more likely to be included in LAD 6.5+, while images mentioning men or LGBTQ+ communities are more likely to be excluded. Also, images with descriptions mentioning Hindu, Buddhist, or Christian communities are more likely to be included in LAD 6.5+, while images with descriptions mentioning Jews or Muslims are more likely to be filtered out. Images mentioning Asian or African communities are more likely to be included in LAD 6.5+ than those mentioning Latinx/e, Caucasian, or European communities. This may suggest that LAP more highly rates images of Asian and African communities than European communities. Alternatively, images containing Caucasians or Europeans may not mention this detail in their captions because whiteness is an unmarked racial category (Brekhus, 1998).
As PMI is a relative measure, one must also consider the absolute number of image descriptions per regex. For example, the PMI of ‘whites?’ and ‘latin[oaxe]s?’ are both -0.95. However, there are 9,617 images in LAD 6.5+ with ‘whites?’ in their description but only 26 images for ther regex ‘latin[oaxe]s?’. Not only does ‘wom[ae]n’ have the highest PMI but the regex also appears in more image descriptions in LAD 6.5+ than any other regex in Figure 2 (15,706). Meanwhile, regexes associated with LGBTQ+ identities are far rarer in LAD 6.5+ image descriptions: ‘gays?’ (102), ‘lesbians?’ (20), ‘transgenders?’ (9), ‘bi-?sexuals?’ (2), and ‘non[ -]?binary’ (0). Understanding the cultural implications of aesthetic evaluation requires considering more than the representation of identity groups (Taylor et al., 2025). Below, we explore aesthetic filtering in greater detail by examining how LAP rates art from various cultures.
Median Score Num Images
Scored $\geq$ 6.5
Num Images
Scored $\geq$ 6
Num Images
Genre Cityscape 5.97 539 2,151 4,602
Portrait 5.90 693 5,616 14,112
Landscape 5.87 505 4,637 13,358
Illustration 5.36 11 140 1,902
Nude Painting 5.45 3 117 1,923
Abstract Painting 4.94 0 0 4,968
Style Realism 5.87 634 4,155 10,733
Romanticism 5.87 371 2,723 7,019
Impressionism 5.89 646 4,898 13,060
Cubism 5.50 12 210 2,235
Pop Art 5.05 2 33 1,483
Abstract Expressionism 4.96 0 7 2782
Artist Antoine Blanchard 6.94 168 170 170
Edouard Cortes 7.09 208 214 214
Pablo Picasso 5.51 1 65 762
Andy Warhol 5.22 0 6 107
Salvador Dalí 5.17 0 20 479
Table 2. A representative selection of the highest and lowest scored genres, styles, and artists in WikiArt. The full results per genre, style, and artist can be downloaded at https://github.com/jtaylor351/laion_aesthetics_audit_data/tree/main
$3.$ Findings: The Metropolitan Museum of Art Dataset